OpenStudy (anonymous):

Given the function f(x) = 0.5(3)x, what is the value of f−1(7)?

OpenStudy (anonymous):

sO far, i have x=0.5(3)y

OpenStudy (mathstudent55):

\(\Large f(x) = (0.5 ) 3^x\) \(\Large 7 = (0.5)3^x\) Solve for x.

OpenStudy (anonymous):

7=1.5^x

OpenStudy (mathstudent55):

No. You can't multiply 0.5 by 3 since 3 is raised to x. What you can do is divide both sides by 0.5. Then take log of both sides.

OpenStudy (mathstudent55):

Look at this example. See why you can't multiply the 0.5 and the 3 together because of the exponent.